11 1.53 points1 PrevicuAnmr My Suppose that a normal model with mean 100 and standard devietion 12 can be epplied to the scores of a reference population on the Wechsler Intelligence Scale for Children (WISC). Children (WISc). Question 1 What ercent of廿is population have wrsc scores below 84.07 (NOTE that requested answer is a percent 9,18 Question 2. What perce nt of this population have WISC scores abave 132.07 Question 3. What nercent of this population have WISC scores between 100 and 116.0 40.82 Question 4. What are the quartes of WISC scores? Submit Answer Save Progress
Answer #1

Solution:
Given: mean = 100 and sd= 12
To find the 1st quartile ,we have to find the corresponding z-values as
P( Z>=z)= 25%
p( Z>=z)= 0.25
P( Z>= - 0.6745)= 0.25
From the standard normal table.
Now using the z-score formula, we can find the value of x as follows,
z=(x-mean)/sd
z*sd =( x- mean)
x= mean +( z* sd)
x= 100+(- 0.6745*12)
=100+(-8.094)
x=91.906
The 1st qurtile is 91.906
Similarly,
To find the 3rd qurtile ,we have to find the corresponding z-values as
P( Z>=z)= 75%
p( Z>=z)= 0.75
P( Z>= 0.6745)= 0.75
From the standard normal table.
Now using the z-score formula, we can find the value of x as follows,
z=(x-mean)/sd
z*sd =( x- mean)
x= mean +( z* sd)
x= 100+(0.6745*12)
=100+(8.094)
x=108.094
The 3rd qurtile is 108.094